Suppose the following demand and supply function:

Qd = 750 – 25P

Qs = -300 + 20 P


      i.         Find the equilibrium price and quantity

    ii.         Find consumer and producer surplus



1
Expert's answer
2021-05-25T10:44:32-0400

(i)"Qd=Qs"

"750-25p=-300+20p"

"750+300=20p+25p"

"1050=45p"

"p=23.33"


"Q=750-25(23.33)"

"=750-583.25"

"Q=166.75"


(ii) consumer/producer surplus"=\frac{1}{2}\times base\times height"

 Ps and Pd are also calculated at points where quantity demanded and supplied are 0. 

"0=750-25P_d"

"25P_d=750"

"Pd=\frac{750}{25}"

"=30"


"o=-300+20P_s"

"20P_s=300"

"P_s=\frac{300}{20}"

"=15"


The consumer and producer surplus are then calculated as follows


consumer surplus

"=\frac{1}{2}\times166.75\times(30-23.33)"

"=\frac{1}{2}\times166.75\times6.67"

"=556.11"


producer surplus

"=\frac{1}{2}\times166.75\times(23.33-15)"

"=\frac{1}{2}\times166.75\times8.33"

"=694.51"